The annual income from an apartment complex is $20220. The annual expense is estimated to be $3435. The apartment complex could be sold for $123686 at the end of 10 years. If your MARR is 10%, how much should you pay for the apartment complex if you were to buy it now? 
Answer #1

ANSWER:

Annual income = $20,220

Annual expense = $3,435

Annual benefit = Annual income - Annual expense = $20,220 - $3,435 = $16,785

Salvage value = $123,686

n = 10 years

i = 10%

Present value = Annual benefit(p/a,i,n) + salvage value(p/f,i,n)

Present value = 16,785(p/a,10%,10) + 123,686(p/f,10%,10)

Present value = 16,785 * 6.145 + 123,686 * 0.3855

Present value = 103,143.83 + 47,680.95

Present value = 150,824.78

So, right now he should pay $150,824.78 for the house.